LgswUnitSx test w/ ASM

We cheked again the PC behavour after 1 night at 40% RH. The BLue and Yellow still look bad:

20161204_145817_pcTransmission.jpg

We could close the loop w/ LGSW only and the max gain was about 0.1, then we had oscillations...
TN AO Gain AO CL TT Res. WF [nm rms] HO Res. WF [nm rms] AO rec DIMM ["] JCL PCL Notes
20161204_174054 0.3,0.2,0.1 T 93.2076 27.1931 20161013_153800 -1 F F LGSW only
20161204_174110 0.3,0.2,0.1 T 88.6853 27.2272 20161013_153800 -1 F F LGSW only
20161204_174120 0.3,0.2,0.1 T 93.4872 27.2478 20161013_153800 -1 F F LGSW only

Checked also loop w/ pyramid as TTS and it worked, also TS was working...

LgswUnitDx calibrations

11:00: setup on DX side finished, we had problems with both telescope authorization and FLAO setup.

We close on LGSW only using 2015/07 rec.
  • Adsec flat: 20161205_argos.sav
  • pupil motors positions: [410; 510], [280; 250], [405; 320]

DX HEXA -3.89 -2.1 0.66 -217 -370
Then we sweep the FoV with the hexapod.
Topic revision: r3 - 28 Apr 2017, MarcoBonaglia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback